A cellular phone company has the following production function for a smart phone: p(x, y) =

50x

2

3 y

1

3 where p is the number of units produced with x units of labor and y units of capital. a)

Find the number of units produced with 125 units of labor and 64 units of capital. b) Find the

marginal productivities (Hints: Partial derivatives). c) Evaluate the marginal productivities at

x = 125 and y = 64.


1
Expert's answer
2021-10-27T14:52:25-0400

Answer:-


"p(x,y)=50x^{2\/3}y^{1\/3}"


a)



"p(125,64)=50(125)^{2\/3}(64)^{1\/3}=5000"

b)



"\dfrac{\partial p}{\partial x}=\dfrac{100}{3}x^{-1/3}y^{1/3}""\dfrac{\partial p}{\partial y}=\dfrac{50}{3}x^{2/3}y^{-2/3}"

c)



"\dfrac{\partial p}{\partial x}(125,64)=\dfrac{100}{3}(125)^{-1/3}(64)^{1/3}=\dfrac{80}{3}""\dfrac{\partial p}{\partial y}(125,64)=\dfrac{50}{3}(125)^{2/3}(64)^{-2/3}=\dfrac{625}{24}"